# Budget-Adaptive Adapter Tuning in Orthogonal Subspaces for Continual Learning in LLMs

### v1

@@ Appendix A Appendix
Table 6 shows details of the 15 datasets we used for our CL experiments, along with their evaluation metrics. Overall, we used datasets from CL benchmark [31], GLUE [32] and SuperGLUE [33] benchmarks, and added IMDB movie reviews dataset [34], following [20]. Table 7 shows details of task orders used for our CL experiments.
@@
Table 6: The details of 15 datasets used in our CL experiments. NLI denotes natural language inference, QA denotes questions and answers task. The first five tasks correspond to the standard CL benchmark, all other tasks are used in long-sequence experiments.
@@
Table 7: Six different task sequence orders utilized in continual learning experiments. Orders 1-3 follow the standard continual learning benchmark as established by previous research, focusing on a more traditional task sequence. Orders 4-6 customized for long-sequence experimentation, encompass 15 tasks each and are structured according to the methodologies outlined in [20].
@@
A.6 Limitations and Further Research Directions.
While our method has demonstrated effectiveness in empirical evaluations, there are a few limitations and potential directions of research to consider. Firstly, although our method does not rely on task identification during inference, it still requires task identification during training. Exploring methods for task-agnostic training would be a valuable future direction. Additionally, while our findings in Figure 2 reveal that models do not completely forget knowledge from previous tasks due to cross-task interference, the underlying mechanisms and potential ways to leverage this phenomenon remain unexplored in our current approach. This represents a promising new perspective and direction for future continual learning research that warrants further investigation.

### v2

@@ A.2 Details About Orthogonality Constraints
Table 6 and  7 show details of the datasets we used for our experiments, along with their evaluation metrics. Overall, we used datasets from the Standard CL benchmark Zhang et al. (2015), GLUE Wang et al. (2019b), and SuperGLUE Wang et al. (2019a) benchmarks, and added the IMDB movie reviews dataset Maas et al. (2011), following Razdaibiedina et al. (2023). Additionally, we incorporated the SuperNI Benchmark Wang et al. (2022a), a benchmark of diverse NLP tasks with expert-written instructions, for more rigorous evaluation of our model. Furthermore, Table 8 shows details of task orders used in our CL experiments.
@@
Table 6: The details of 15 classification datasets in the Long Sequence Benchmark. NLI denotes natural language inference, QA denotes questions and answers task. First five tasks correspond to the standard CL benchmark, all other tasks are used in long-sequence experiments.
@@
Table 7: The details of 15 datasets in the SuperNI Benchmark.
Dataset name Task Metric
task639_multi_woz_user_utterance_generation dialogue generation Rouge-L
task1590_diplomacy_text_generation dialogue generation Rouge-L
task1729_personachat_generate_next dialogue generation Rouge-L
task181_outcome_extraction information extraction Rouge-L
task748_glucose_reverse_cause_event_detection information extraction Rouge-L
task1510_evalution_relation_extraction information extraction Rouge-L
task002_quoref_answer_generation question answering Rouge-L
task073_commonsenseqa_answer_generation question answering Rouge-L
task591_sciq_answer_generation question answering Rouge-L
task511_reddit_tifu_long_text_summarization summarization Rouge-L
task1290_xsum_summarization summarization Rouge-L
task1572_samsum_summary summarization Rouge-L
task363_sst2_polarity_classification sentiment analysis accuracy
task875_emotion_classification sentiment analysis accuracy
task1687_sentiment140_classification sentiment analysis accuracy
Table 8: Eight different task sequence orders utilized in our experiments. Orders 1-3 follow the standard continual learning benchmark as established by previous research, focusing on a more traditional task sequence. Orders 4-6 are customized for long-sequence experimentation, encompassing 15 tasks each and are structured according to the methodologies outlined in Razdaibiedina et al. (2023). Orders 7-8 correspond to the SuperNI benchmark.
@@
7 task1572 $\rightarrow$ task363 $\rightarrow$ task1290 $\rightarrow$ task181 $\rightarrow$ task002 $\rightarrow$
task1510 $\rightarrow$ task639 $\rightarrow$ task1729 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$
task748 $\rightarrow$ task511 $\rightarrow$ task591 $\rightarrow$ task1687 $\rightarrow$ task875
8 task748 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$ task639 $\rightarrow$ task1572 $\rightarrow$
task1687 $\rightarrow$ task591 $\rightarrow$ task363 $\rightarrow$ task1510 $\rightarrow$ task1729 $\rightarrow$
task181 $\rightarrow$ task511 $\rightarrow$ task002 $\rightarrow$ task1290 $\rightarrow$ task875
A.4 Implementation Detail.
All our experiments involving T5 models were performed on a server outfitted with four NVIDIA GeForce RTX 3090 GPUs, utilizing the DeepSpeed repository for implementation. Following previous studies de Masson d’Autume et al. (2019); Rao et al. (2019), for CL experiments, for each dataset we use the available validation set as a test set (since test data is not available) and hold out 500 samples from the train set to construct the validation set. For every sequence of tasks across different orders, we trained the models for one epoch using a batch size of 32 (8 per GPU), a dropout rate of 0.1, and no weight decay. Across all experiments, we primarily used Adapter modules with a bottleneck dimension of 16, and applied a sparsification threshold chosen from $\{1e\text{-}3,1e\text{-}4,1e\text{-}5\}$. The learning rate was selected from $\{5e\text{-}3,3e\text{-}3,1e\text{-}3,5e\text{-}4\}$ depending on task characteristics. We applied an orthogonality regularization on the Adapter’s upsampling matrix with a coefficient $\lambda_{orth}\in\{0.5,1,5\}$, and used an additional coefficient $\lambda_{2}\in\{0,0.1,0.5\}$ to scale the associated L2 loss term. For T5-Large, we report the average accuracy (AA) for Orders 1-3 in Table 9, as the primary metric for evaluating the effect of $\lambda_{\text{orth}}$. Unless otherwise specified, the reported main results use the $\lambda_{\text{orth}}$ selected by the best validation AA averaged across orders. This protocol balances knowledge retention and sparsity, and yields stable performance across long sequences with distribution shift. To ensure experimental comparability and fair result comparison, we maintain consistency with O-LoRA Wang et al. (2023a) by adopting instruction tuning as the training paradigm across all experiments for both our method and other baselines, as shown in Table 10. This approach offers dual advantages: it incorporates human expertise for efficient learning while enabling models to better capture underlying principles through explicit guidance, thereby enhancing generalization capabilities. The consistent instruction-based framework allows for direct performance comparisons while leveraging the benefits of natural language supervision.
Table 9: Effect of the orthogonality coefficient $\lambda_{\text{orth}}$ on T5-large.
$\lambda_{\text{orth}}$ Order 1 Order 2 Order 3 Average
0.5 74.3 75.3 74.5 74.7
1 73.4 74.8 75.1 74.4
2 73.5 74.9 73.2 73.9
3 74.2 73.9 74.1 74.1
4 74.7 76.4 73.6 74.9
5 75.8 75.6 74.6 75.3
Table 10: Instructions for different tasks.
Task Prompts
NLI What is the logical relationship between the ”sentence 1” and the ”sentence 2”?
Choose one from the option.
QQP Whether the ”first sentence” and the ”second sentence” have the same meaning?
Choose one from the option.
Choose one from the option.
or false? Choose one from the option.
in both sentence? Choose one from the option.
A.5 Heterogeneous Budget Requirements Across Tasks and Layers.
As discussed in Section 4.2, we extend our analysis to investigate budget allocation in the context of continual learning. Here, we further present results under other task orders, as illustrated in Figure 4 and 5. These findings corroborate our analysis in Section 4.2: (a) The relationship between performance and parameter budget does not follow constant rules but rather necessitates case-specific consideration. (b) Within continual learning scenarios, the first task primarily focuses on acquiring capabilities built upon the pretrained model, whereas subsequent tasks must additionally preserve knowledge from preceding tasks, thus requiring more nuanced fine-tuning. This further substantiates our analysis that different tasks in the CL scenarios require varying budgets, and that allocating budgets according to training sequence and task characteristics is both necessary and justified.
A.6 Occurrence and Mitigation of Catastrophic forgetting.
To further validate the effectiveness and consistency of orthogonal parameter subspace constraints, we conduct sequential training following Order-2 and Order-3, with results illustrated in Figure 6 and 7, respectively. Consistent with the results reported in Section 4.3, we observe severe catastrophic forgetting in the absence of orthogonal constraints, especially for earlier tasks. In contrast, models trained with orthogonal parameter subspace constraints are able to preserve performance across all tasks to a much greater extent. Notably, although the specific tasks affected most by forgetting vary depending on the task order, the general trend holds: orthogonal constraints provide consistent mitigation of forgetting regardless of task permutation. These results reinforce our earlier findings and highlight the robustness of orthogonal subspace regularization as a general mechanism for alleviating forgetting in continual learning scenarios.
A.7 Compatibility with Replay.
In this appendix, we discuss the compatibility of our method, OA-Adapter, with replay-based strategies. We compare this hybrid with SAPT (Zhao et al., 2024), a strong baseline in the replay-based continual learning setting.
Our experiments, conducted on the SuperNI Benchmark, show that the OA-Adapter+Replay hybrid consistently achieves additional gains in performance while maintaining low forgetting compared to the original OA-Adapter. Specifically, the average performance of our method is 48.3, outperforming SAPT+Replay’s 44.2 on the SuperNI dataset. This indicates that our method can effectively complement replay strategies and improve performance.
Although OA-Adapter+Replay achieves lower performance than SAPT+ARM (50.8), we note that our approach still demonstrates better compatibility with replay-based strategies, offering a more efficient and scalable solution.
However, it is important to note that replay-based methods come with significant overhead, either requiring the storage of large amounts of task data or generating synthetic data for retraining. Both approaches introduce challenges, including storage limitations and potential privacy concerns due to data retention. Additionally, replay-based methods often incur increased computational costs, as they necessitate continuous retraining with stored or generated data.
Given these challenges, while our method demonstrates compatibility with replay-based strategies and provides notable improvements, we chose not to incorporate replay into our final approach due to the associated resource-intensive nature and practical limitations. Instead, we focus on efficient, memory-conscious strategies that can be more easily scaled in real-world applications without the need for large-scale data storage or generation.
A.8 Compute and Memory Cost Analysis
Training cost and GPU memory.
We benchmark the training cost of OA-Adapter against O-LoRA under Task Order 1 on the Standard CL benchmark. OA-Adapter demonstrated slightly lower GPU memory usage, averaging 68.9% compared to 70.4% for O-LoRA. OA-Adapter achieved a significantly faster average training time per step: 0.357 seconds versus 0.582 seconds for O-LoRA, suggesting a clear advantage in training efficiency.
Overhead from orthogonality regularization.
We further isolate the per-step time consumed by orthogonality computations for OA-Adapter (bottleneck dimension 24) on the Standard CL benchmark, as shown in Table 11. The per-step overhead grows with the number of previously seen tasks, resulting in an average of 0.0179 s per step across the first four tasks, i.e., a 4.6% relative overhead compared to the total per-step time of 0.357 s. As expected from the pairwise constraints with prior tasks, this cost grows approximately linearly with the task sequence length; extrapolating to a 15-task sequence yields an estimated proportion of 14.7% of total training time devoted to orthogonality operations.
Table 11: Per-step time spent on orthogonality computations for OA-Adapter.
Task 1 2 3 4 avg
Time (s) 0.0061 0.0130 0.0199 0.0269 0.0179